# Linearly Multiplexed Photon Number Resolving Single-photon Detectors Array

### v1

@@ Abstract
I Introduction
@@
Photon Number Resolving Detectors can be realized with different technologies, each one with its characteristics, operating principles, and limitations. For instance, schemes with photomultiplier tubes [12] and Geiger-mode avalanche photodiodes (GAPD) [13] have shown photon number resolution with few photons. Visible light photon counters [14] have a high dark count rate and slow operational rate, while transition edge sensors (TES) can discriminate the number of photons with high efficiency (98%) [15] and low noise at telecommunication wavelengths, but their performance is limited by the low count rates and high time jitter [16, 17].
Superconducting Nanowire single photon detector (SNSPD) [18] is considered one of the most promising platform for overcoming the limitations of the detector platforms mentioned above, thanks to its high efficiency [19], negligible dark count rate [20], jitter time in the picosecond region, [21] and short recovery time [22]. However, single SNSPD detectors do not have the intrinsic photon number resolution capability like TES detectors. Here we show how this limit can be addressed with a multiplexed scheme on an optical waveguide to enable photon number resolution and optimized fidelity.
Integration of SNSPD detectors on optical waveguide chips has been demonstrated, showing how these devices maintain their outstanding properties for a diverse range of materials [23]. Integration with complex photonic waveguide networks has shown the possibility of creating a combined platform which contains both capabilities of SNSPDS and PNRDs with great scalability potential [24, 25, 26] and applications in quantum communication [27], characterization of quantum photon statistics [25] and quantum sensing [28].
@@
The manuscript is structured as follows: In Sect. II we will discuss a general approach for the geometry of an array of SNSPDs on a linear waveguide and calculate the relevant merit factors in the case of ideal waveguide and detectors. Then, non-idealities such as propagation losses and detector dark counts will be discussed and estimated. In Sect. III an actual design for the implementation of the waveguide PNRD will be proposed and discussed. Simulations will be used to calculate the relevant parameters for the proposed device so that the final merit factors can be calculated using the theory developed in Sect. II. In Section IV we will summarize the results and draw our conclusions.
II A multiplexed SNSPDs on a single waveguide
II-A Lossless waveguide and detector
@@
The fidelity of Eq. 6 is shown in Fig. 2 for an array of $N=120$ elements, and it represents the ideal case in which we considered only dichotomous events of photons being either absorbed or not.
@@
II-B Propagation losses and fidelity
@@
II-C Detector with dark counts
@@
III Design of the detector array on a strip-loaded waveguide on TFLN
The relations for the fidelity derived in the previous sections can be applied to any material platform for either the waveguide and the nanowire detectors. Now we consider a device fabricated on a lithium niobate on insulator (LNOI) substrate loaded with a Silicon Nitride (Si3N4 or SiN) waveguide strip and coupled to a series of Niobium Nitride (NbN) SNSPDs, with the geometry sketched in Fig. 1. Figure 4 shows a schematic and cross section of this configuration.
Figure 4: Cross section of the photon-number resolving detector: starting from the bottom there are a crystalline silicon ($c$-Si) layer, a silicon oxide (SiO2) layer, a lithium niobate (LiNbO3) layer, a silicon nitride (SiNx) waveguide surrounded by an SiO2 layer, and finally the niobium nitride (NbN) layer. The detailed dimensions are reported in Tab.I.
The device is made on a the substrate with a $300$ nm lithium niobate layer on a 3 $\mu$m silicon dioxide (SiO2) insulator. On the LNOI film a ridge in SiN is patterned with a nominal a width of 1.2 $\mu$m and a thickness of 200 nm, defining the waveguide channel. The superconducting nanowire are patterned on top of the SiN ridge with a fixed thickness of $6$ nm and a width of 80 nm. This structure is simulate using the finite elements method (FEM) with the parameters summarized in Tab. I.
@@ III Design of the detector array on a strip-loaded waveguide on TFLN
Refractive index $n$ 1.44 $\begin{gathered}2.21(o)\\ 2.14(e)\end{gathered}$ 1.88 $5.825.23+i$ 5.82
@@
TABLE I: List of parameters used for the FEM simulations
III-A Straight and curved propagation
Initially, to estimate the propagation losses in passive structures, and determine $\alpha_{w}$, FEM simulations were conducted for strip-loaded waveguides, excluding the NbN layer, for both straight and curved configurations. Additionally, in order to assess the stability of the confined mode with respect to leakages into slab modes, simulations excluding the SiN layer were conducted and compared to the strip-loaded ones [33].
For the first part of simulations, we performed a parametric sweep varying both the SiN waveguide width and height to take into account different fabrication scenarios. The waveguide height ranged from 180 to 220 nm, while the width varied between 1 $\mathrm{\mu m}$ and 1.4 $\mathrm{\mu m}$. For the parameters specified in Tab. I we obtained a single mode waveguide, with an effective refractive index $n_{\text{eff}}=1.8066$.
@@ III-A Straight and curved propagation
III-B Tuning SPDs efficiency
@@
Despite its flexibility, this configuration poses challenges in bending the nanowires over the waveguide. Indeed, curved nanowires often lead to current crowding issues [36], making straight sections practically ineffective. To mitigate this occurrence, widening the nanowire in curved sections is a potential solution, but it introduces absorptive dead zones atop the waveguide, significantly impacting the fidelity.
@@
The values obtained for each $i$ between 1 and 33 are reported in Tab. II.
@@
TABLE II: List of $u_{i}$ meanders of each SPD
Using these data and the methodologies discussed in Sec. II and the worst dark counts scenario in [26] we evaluate the performance of a PNRD composed by $N=33$ meander SPDs in different regimes: ideal, lossy and lossy plus DC. We then compare these results to the non-discretized ideal scenario by examining the fidelity ratio ${F_{\text{disc}}}/{F_{\text{ideal}}}$, as shown in Fig. 8: remarkably, the introduction of losses did not change significantly the performance of such PNRD, which the introduction of dark counts did instead.

### v2

@@ Abstract
[…]
[IFN] organization=Istituto di Fotonica e Nanotecnologie, CNR-IFN, addressline=Via del Fosso del Cavaliere 100, postcode=00133, state=Roma, country=Italy
[…]
[TIFPA] organization=INFN TIFPA, Trento Institute for Fundamental Physics and Applications, addressline=Via Sommarive 14, city=Povo, postcode=38123, state=Trento, country=Italy
1 Introduction
@@
Photon Number Resolving Detectors can be realized with different technologies, each one with its characteristics, operating principles, and limitations. For instance, schemes with photomultiplier tubes [12] and Geiger-mode avalanche photodiodes (GAPDs) [13] have shown photon number resolution with few photons. Visible light photon counters [14] have a high dark count rate and slow operational rate, while transition edge sensors (TES) can discriminate the number of photons with high efficiency (98%) [15] and low noise at telecommunication wavelengths, but their performance is limited by the low count rates and high time jitter [16, 17].
Superconducting Nanowire single photon detector (SNSPD) [18] is considered one of the most promising platform for overcoming the limitations of the detector platforms mentioned above, thanks to its high efficiency [19], negligible dark count rate [20], jitter time in the picosecond region, [21] and short recovery time [22]. However, single SNSPD detectors do not have intrinsic photon number resolution capability like TES detectors. Here we show how this limit can be addressed with a multiplexed scheme on an optical waveguide to enable photon number resolution and optimized fidelity.
Integration of SNSPD detectors on optical waveguide chips has been demonstrated, showing how these devices maintain their outstanding properties for a diverse range of materials [23]. Integration with complex photonic waveguide networks has shown the possibility of creating a combined platform which contains both capabilities of SNSPDS and PNRDs with great scalability potential [24, 25, 26, 27, 28] and applications in quantum communication [29], characterization of quantum photon statistics [25] and quantum sensing [30].
@@
The manuscript is structured as follows: In Sect. 2 we will discuss a general approach for the geometry of an array of SNSPDs on a linear waveguide and calculate the relevant merit factors in the case of ideal waveguide and detectors. Then, non-idealities such as propagation losses and detector dark counts will be discussed and estimated. In Sect. 3 an actual design for the implementation of the waveguide PNRD will be proposed and discussed. Simulations will be used to calculate the relevant parameters for the proposed device so that the final merit factors can be calculated using the theory developed in Sect. 2. In Section IV we will summarize the results and draw our conclusions.
2 A multiplexed SNSPDs on a single waveguide
2.1 Lossless waveguide and detector
@@
It is worth to note that the same result is obtained in [34] by starting from the same assumptions. The fidelity of Eq. 6 is shown in Fig. 2 for an array of $N=120$ elements, and it represents the ideal case in which we considered only dichotomous events of photons being either absorbed or not.
@@
2.2 Propagation losses and fidelity
@@
2.3 Detector with dark counts
@@
3 Design of the detector array on a strip-loaded waveguide on TFLN
The relations for the fidelity derived in the previous sections can be applied to any material platform for either the waveguide and the nanowire detectors. Now we consider a device fabricated on a lithium niobate on insulator (LNOI) substrate loaded with a Silicon Nitride (Si3N4 or SiN) waveguide strip and coupled to a series of Niobium Nitride (NbN) SNSPDs, with the geometry sketched in Fig. 4a. The PNRD is fabricated on a LNOI substrate made by a $300$ nm lithium niobate layer on a 3 $\mu$m silicon dioxide (SiO2) insulator. On the LNOI film, a ridge in SiN is patterned with a nominal a width of 1.2 $\mu$m and a thickness of 200 nm, defining the waveguide channel. The SNSPDs are then patterned on top of the SiN ridge with a fixed thickness of $6$ nm and a width of 80 nm. Such series of SNSPDs can be realized on top of the strip-loaded LNOI photonic circuit thanks to the following fabrication steps. A SiO2 cladding layer is deposited by Plasma-Enhanced Chemical Vapor Deposition, and a planarization step is performed by Chemical Mechanical Polishing (CMP). The NbN superconducting film is deposited using DC Magnetron Sputtering, and the different nanowire patterns are realized by Electron Beam Lithography followed by dry etching. AuPd resistors, placed in parallel to each nanowire element, and Au contact pads are fabricated via lift-off. Such fabrication procedure is the same employed in [36], except for the CMP step that allows to position both the contact pads and AuPd parallel resistors on the cladding layer. In addition, such step grants optimal surface roughness prior the SNSPD fabrication.
Such a structure is simulated using the finite elements method (FEM) with the parameters summarized in Tab. 1.
Figure 4: Cross sections of the photon-number resolving detector: starting from the bottom there are a crystalline silicon ($c$-Si) layer, a silicon oxide (SiO2) layer, a lithium niobate (LiNbO3) layer, a silicon nitride (SiNx) waveguide surrounded by an SiO2 layer, and finally the niobium nitride (NbN) layer. The detailed dimensions are reported in Tab.1.
@@ 3 Design of the detector array on a strip-loaded waveguide on TFLN
Refractive index $n$ 1.44 $\begin{gathered}2.21(o)\\ 2.14(e)\end{gathered}$ 1.88 $\begin{gathered}5.23+\\ i5.82\end{gathered}$
@@
Table 1: List of parameters used for the FEM simulations. Both SiO2 and SiN refractive index are obtained by Variable Angle Spectroscopic Ellipsometry measurements on fabricated samples. LiNbO3 and NbN refractive index are respectively reported in [37] and [38].
3.1 Straight and curved propagation
Initially, to estimate the propagation losses in passive structures, and determine $\alpha_{w}$, FEM simulations were conducted for strip-loaded waveguides, excluding the NbN layer, for both straight and curved configurations.
For the first part of simulations, we performed a parametric sweep varying both the SiN waveguide width and height to take into account different fabrication scenarios. The waveguide height ranged from 180 to 220 nm, while the width varied between 1 $\mathrm{\mu m}$ and 1.4 $\mathrm{\mu m}$. For a thickness of 200 nm and a width of 1.2 $\mu$m, we obtained a single mode waveguide with an effective refractive index $n_{\text{eff}}=1.8066$.
@@ 3.1 Straight and curved propagation
3.2 Tuning SPDs efficiency
@@
Despite its flexibility, this configuration poses challenges in bending the nanowires over the waveguide. Indeed, curved nanowires often lead to current crowding issues [42], making straight sections practically ineffective. To mitigate this occurrence, widening the nanowire in curved sections is a potential solution, but it introduces absorptive dead zones atop the waveguide, as noticeable in Fig. 4b, significantly impacting the fidelity.
@@
The values obtained for each $i$ between 1 and 33 are reported in Tab. 2.
@@
Table 2: List of $u_{i}$ meanders of each SPD
Using these data and the methodologies discussed in Sec. 2 and the worst dark counts scenario in [26] we evaluate the performance of a PNRD composed by $N=33$ meander SPDs in different regimes: ideal, lossy and lossy plus DC. We then compare these results to the non-discretized ideal scenario by examining the fidelity ratio ${F_{\text{disc}}}/{F_{\text{ideal}}}$, as shown in Fig. 8: remarkably, the introduction of losses did not change significantly the performance of such PNRD, which the introduction of dark counts did instead.